OpenStudy (anonymous):

Use the following equation to find the value of y when x = 3 8x - 10y = 39

OpenStudy (bibby):

plug in x =3 combine like terms solve for y

OpenStudy (anonymous):

8x - 10y = 39 8(3)- 10y = 39 24-10y=39 -24 -24 -------------- -10y/-10=15/-10 y=3/-2

OpenStudy (anonymous):

@bibby

OpenStudy (vickyisthesmartone123):

Use mathway.com it gives you correct answers!!!:)

OpenStudy (bibby):

8x - 10y = 39 x=3 8(3)-10y=39 24-10y=39 -10y=15 y=-15/10 y=-3/2

OpenStudy (bibby):

correct